OpenStudy (anonymous):

Which is a factor of 4x^2y + 24x^2 - 24y - 144? a. 3 b. y + 4 c. x^2 - 4 d. y + 6

OpenStudy (anonymous):

= 4(x^2y + 6x^2 - 6y - 36) = 4(y(x^2 - 6) + 6(x^2 - 6)) = 4(x^2-6)"(y + 6)" D